OpenStudy (anonymous):

order the quadratic functions in order from widest to narrowest graph y=5x^2,y=x2,y=0.75x^2

OpenStudy (welshfella):

the 5 in front of the x^2 stretches the graph of y^2 parallel to the y-axis so it will be narrower thsn y = x^2

OpenStudy (welshfella):

the 0.75 before the x^2 will spread the graph of y=x^2 out

OpenStudy (welshfella):

- widen it

OpenStudy (anonymous):

so it would be y=x^2,y0.75x^2,y=5x^2

OpenStudy (welshfella):

no - widest first

OpenStudy (anonymous):

y=0.75x^2,y=x^2,y=5x^2

OpenStudy (welshfella):

thats it